OpenStudy (anonymous):

What is the x-intercept and y-intercept of the graph of 10x - 5y = 20?

OpenStudy (anonymous):

All answerers will recive medals (:

OpenStudy (anonymous):

x intercept (2,0) y -------->(0,-4)

OpenStudy (anonymous):

x-intercept | 2 y-intercept | -4 slope | 2

OpenStudy (anonymous):

X intercept Y intercept 10x-5y=20 10x-5y=20 10x-0y=20 0x-5y=20 10x=20 5y=20 /10=/10 /5=/5 x=2 y=4

OpenStudy (anonymous):

x=2, y=-4